A website reports that 75​% of its users are from outside a certain country and that 40​% of its users log on every day. Suppose that 11​% of its users are users from the country who log on every day. Make a probability table. Why is a table better than a tree​ here?

Complete the probability table below.

From the Country

Not from the Country

Total

Log on Every Day

Do Not Log on Every Day

Total

1.00

​(Type integers or​ decimals.)
Answer #1

since
P(not from the country) =0.75

P(from the country) =1-0.75 =0.25

P(log on every day) =0.40

P(do not log on every day) =1-0.40 =0.6

below is probability table:

from the country Not from the country total
every day 0.11 0.29 0.4
do not log on every day 0.14 0.46 0.6
total 0.25 0.75 1